OpenStudy (anonymous):

what are the slopes of two lines that are prallel and perpendicular to line 4x-8y=7

OpenStudy (anonymous):

4/8

OpenStudy (anonymous):

sorry 8/4

OpenStudy (anonymous):

perpendicular is the negative reciprocal of the slope of the line. parallel have same slope

OpenStudy (unklerhaukus):

4x-8y=7 -8y=-4x+7 y=-x/2-7/8 slope m = -1/2 parallel lines will have this slope, perpendicular lines will have slope of 2

OpenStudy (anonymous):

thanks!